We roll a pair of dice once and are given that the two numbers that occur are not the same. Compute the following probabilities:

• i) the sum is 7

• ii) the sum is 4

• iii) the sum is 12



1
Expert's answer
2022-05-18T17:46:19-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & - & 3 & 4 & 5 & 6 & 7 \\
 \hdashline
 2 & 3 & - & 5 & 6 & 7 & 8 \\
 \hdashline
 3 & 4 & 5 & - & 7 & 8 & 9 \\
 \hdashline
 4 & 5 & 6 & 7 & - & 9 & 10 \\
 \hdashline
 5 & 6 & 7 & 8 & 9 & - & 11 \\
 \hdashline
 6& 7 & 8 & 9 & 10 & 11 & - \\
 \hdashline
\end{array}"

i)


"P(sum=7)=\dfrac{6}{30}=\dfrac{1}{5}"


ii)


"P(sum=4)=\dfrac{2}{30}=\dfrac{1}{15}"


iii)


"P(sum=12)=0"